# P. Guiraud, M. Hernández, A. Meyroneinc and A. Nogueira

### v1

@@ 2.2 Main results
1) As shown in the right panel of Figure 4, for $\rho=p/q$ rational the values of $\delta$ and $a$ satisfying (20) for $\alpha\in(l/q,(l+1)/q)$, where $l\in\{0,\dots,q-1\}$, also satisfy (20) for $\alpha^{\prime}=l/q$. It follows that (23) holds with $\phi_{\delta,\rho,\alpha}$ and $\phi_{\delta,\rho,\alpha^{\prime}}$. As $\phi_{\delta,\rho,\alpha}([0,1))$ is the union of two periodic orbits and $\phi_{\delta,\rho,\alpha^{\prime}}([0,1))$ is a unique periodic orbit, we deduce that the inclusion (23) can indeed be strict.
2) The function $\phi_{\delta,\rho,\alpha}$ allows to parametrize a large variety of Cantor sets. Indeed, choose any $\rho\in(0,1)\setminus\mathbb{Q}$ any $\alpha\in[0,1]$. Then the closure of $\phi_{\delta,\rho,\alpha}([0,1))$ is a Cantor set, whose extremities are 00 and $1$ if $\delta=\delta(\rho,\alpha)$ and $\{\alpha+k\rho\}\neq 0$ for all $k\geq 1$, see Lemma 6.2. Its points are given by
$\phi_{\delta,\rho,\alpha}(y)=\frac{\delta(\rho,\alpha)}{1-\lambda}+\frac{1}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}\left((1-\lambda)\lfloor y-k\rho\rfloor+d\theta_{\alpha}(\{y-k\rho\})\right)\quad\forall y\in\mathbb{R}$ (24)
and the left-sided limits of $\phi_{\delta,\rho,\alpha}$ at its discontinuity points. The closure of $\phi_{\delta,\rho,1}([0,1))$ is the Cantor set $C_{1}$ corresponding to the attractor of the map $x\mapsto\lambda x+\delta\pmod{1}$. In [6], it has been shown that $C_{1}\setminus\{0,1\}$ contains only transcendental numbers if $\lambda=1/b$, where $b\geq 2$ is an integer. It would be interesting to study the persistence of such a property for the Cantor set obtained when $\alpha\neq 1$ as a function of the values of $d$ and $\alpha$.
[…]
@@ 3 Construction and general properties of $\phi$
is the lift of the rotation of angle $\rho$. In this section, we first show that the function $\phi=\phi_{\delta,\rho,\alpha}$ given in (19) is a natural candidate and then we establish conditions which ensure (25).
@@ 3.3 Monotonicity properties of $\phi$
We have proved that $\psi_{\alpha}$ is non-decreasing and item 1)1)1 ). Finally, if $z\notin\mathbb{Z}\cup(\mathbb{Z}+\alpha)$ then, the integer part and the fractional part are continuous at $z$ and $\theta_{\alpha}$ is continuous at $\{z\}$, which implies that $\psi_{\alpha}(z)=\psi_{\alpha}(z^{-})$. Suppose $z\in\mathbb{Z}\cup(\mathbb{Z}+\alpha)$. If $\alpha\in(0,1)$, then either $z\in\mathbb{Z}$ and $\psi_{\alpha}(z)-\psi_{\alpha}(z^{-})=1-\lambda-d$ or $z\in\mathbb{Z}+\alpha$ and $\psi_{\alpha}(z)-\psi_{\alpha}(z^{-})=d$. If $\alpha\in\{0,1\}$, then $z\in\mathbb{Z}\cap(\mathbb{Z}+\alpha)$ and $\psi_{\alpha}(z)-\psi_{\alpha}(z^{-})=1-\lambda$. This ends the proof of (31). ∎
@@
The continuity properties of $\phi$ follow from those of $\psi_{\alpha}$ stated in Lemma 3.5. Now, if $y^{\prime}>y\in\mathbb{R}$, then by 1)1)1 ) of Lemma 3.5
@@
which proves that $\phi$ is a non-decreasing function. Also, by (32), $\phi(y^{\prime})-\phi(y)=0$ if and only if $\psi_{\alpha}(y^{\prime}-k\rho)-\psi_{\alpha}(y-k\rho)=0$ for every $k\geq 1$. By 1)1)1 ) of Lemma 3.5, this can hold only if
@@
If $\rho\in(0,1)\setminus\mathbb{Q}$, then the sequence ($(\{y^{\prime}-k\rho\})_{k\in\mathbb{N}}$ is dense in $[0,1]$. Therefore, there exists $k\in\mathbb{N}$ such that $0\leq\{y^{\prime}-k\rho\}<y^{\prime}-y=y^{\prime}-k\rho-(y-k\rho)$ and
@@
Let $\alpha\in[0,1]$. If $\rho\in(0,1)\cap\mathbb{Q}$, then $\phi([0,1))\subset[0,1)$ if and only if $0\in(\phi(0^{-}),\phi(0)]$. If $\rho\in(0,1)\setminus\mathbb{Q}$, then the same holds replacing $(\phi(0^{-}),\phi(0)]$ with $[\phi(0^{-}),\phi(0)]$.
@@
For $\rho\in(0,1)\setminus\mathbb{Q}$, 1) and 2) hold replacing $(\phi(\alpha^{-}),\phi(\alpha)]$ with $[\phi(\alpha^{-}),\phi(\alpha)]$.
@@
1)1)1 ) Let $\alpha\in(0,1)$. Suppose that $\theta_{\alpha}(y)=\theta_{a}(\phi(y))$ for all $y\in[0,1)$. So, for every $y\in[0,1)$ we have, $\phi(y)<a$ if and only $y<\alpha$. This implies that $\phi(\alpha^{-})\leq a\leq\phi(\alpha)$. In the case $\rho\in(0,1)\cap\mathbb{Q}$, by Lemma 3.8, there exists $y\in[0,\alpha)$ such that $\phi(y)=\phi(\alpha^{-})$ and therefore $\phi(\alpha^{-})<a$.
@@
2)2)2 ) Let $\alpha=0$. If $\phi([0,1))\subset[0,1)$, by Lemma 3.9 we have $\phi(\alpha^{-})=\phi(0^{-})<0\leq a$ if $\rho\in(0,1)\cap\mathbb{Q}$ and $\phi(\alpha^{-})=\phi(0^{-})\leq 0\leq a$ if $\rho\in(0,1)\setminus\mathbb{Q}$. Suppose $1=\theta_{0}(y)=\theta_{a}(\phi(y))$ for every $y\in[0,1)$. Then, $\phi(\alpha)=\phi(0)\geq a$ and it follows that $a\in[\phi(\alpha^{-}),\phi(\alpha)]$ with left open interval if $\rho\in(0,1)\cap\mathbb{Q}$. Conversely, if $a\leq\phi(\alpha)=\phi(0)$, then, by monotonicity of $\phi$, $a\leq\phi(y)$ and $1=\theta_{a}(\phi(y))=\theta_{0}(y)$ for every $y\in[0,1)$.
@@
$=\frac{1}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}(\psi_{\alpha}(\alpha-k\rho)-\psi_{\alpha}(\alpha-k\rho)^{-})$
$=\frac{1}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}\left((1-\lambda-d)\mathbb{1}_{\mathbb{Z}}(\alpha-k\rho)+d\mathbb{1}_{\mathbb{Z}+\alpha}(\alpha+k\rho)\right).$
@@
First, $rp/q\in\mathbb{Z}$ only if $r=0$, since $p$ and $q$ are co-prime. Second, the set $D_{\rho,\alpha}$ is either empty or contains a unique element $r_{\alpha}$. Indeed, if there exist $r_{0}$ and $r_{1}$ in this set, then $|r_{1}-r_{0}|/q\in\mathbb{Z}$ because $p$ and $q$ are co-prime, and it follows that $r_{0}=r_{1}$ since $10\leq|r_{1}-r_{0}|/q<$ 1. Also, the set $\{r\in\{1,\dots,q-1\}:-rp/q\in\mathbb{Z}+\alpha\}$ is empty if and only if $D_{\rho,\alpha}$ is empty and, if it is not empty, then it contains a unique element $r^{\prime}_{\alpha}=-(r_{\alpha}-q)$. It follows that,
@@
The following two lemmata give the the monotonicity properties of $\delta(\rho,\alpha)$ with $\rho$ and allow to show item 3) of Theorem 2.3.
Lemma 4.7.
Also, for any $\rho\in(0,1)$ we have $[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda-d,1)$.
Proof.
As $\psi_{1-\alpha}$ is right continuous, we have
On the other hand,
Lemma 4.8.
Let $\rho\in(0,1)$, $\alpha\in[0,1]$ and suppose $\delta\in[\delta(\rho^{-},\alpha),\delta(\rho^{-},\alpha)]$. Then, $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]\neq\emptyset$ and $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,1]$ if $\alpha\in(0,1)$ or $\rho\notin\mathbb{Q}$.
Proof.
As $\delta\in[\delta(\rho^{-},\alpha),\delta(\rho^{-},\alpha)]$, we deduce from (38) that $\phi(0^{-})\leq 0\leq\phi(0)$, which together with (28) implies $\phi(1^{-})\leq 1\leq\phi(1)$. Since $\phi$ is non-decreasing, we have $0\leq\phi(\alpha)$ and $\phi(\alpha^{-})\leq 1$ for every $\alpha\in[0,1]$. From (40) we obtain that $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]\neq\emptyset$. Now, if $\alpha\in(0,1)$ then $0\leq\phi(\alpha^{-})\leq\phi(\alpha)\leq 1$. If $\alpha\in\{0,1\}$ and $\rho\notin\mathbb{Q}$, then $\phi$ is continuous in $\{0,1\}$ and $\phi(\alpha^{-})=\phi(\alpha)\in\{0,1\}$. We conclude from (38) that $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,1]$ if $\alpha\in(0,1)$ or $\rho\notin\mathbb{Q}$. ∎
For the sake of completeness, we give in the following lemma the expressions of $\delta(\rho,\alpha)$ and $a(\delta,\rho,\alpha)$ for $\rho\in\mathbb{Q}$.
@@
Let $\rho\in(0,1)$ and $\alpha\in[0,1]$. The items 1) and 2) of Theorem 2.3 follow from Lemma 4.5 and Lemma 4.6. The item 3) is deduced from Lemma 4.7 and Lemma 4.8. Now, suppose that $\delta\in(1-\lambda-d,1)$ and $a\in[0,1]$ satisfy
@@
Also, $\phi([0,1))\subset[0,1)$ by Proposition 3.4, which implies that $\left\lfloor\phi(y)\right\rfloor=\left\lfloor y\right\rfloor$ for all $y\in\mathbb{R}$ by 2)2)2 ) of Lemma 3.3. Let $x=\phi(0)$, then
@@
To end the proof of Theorem 2.3, let us suppose that $\rho\in\mathbb{Q}$ and that $(\delta,a)$ satisfies (12) with $\delta=\delta(\rho,\alpha)$ or $a=a(\delta,\rho^{+},\alpha)$. Using continuity arguments of the rotation number given in [25] and adapted to our two parameters context in Appendix A.1, let us show that the rotation number of $F$ is $\rho$. As $\rho$ is a rational number, by Lemma 4.6, the intervals $[\delta(\rho^{-},\alpha),\delta(\rho,\alpha))$ and $(a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]$ are non-empty. So, let $(\delta_{n})_{n\in\mathbb{N}}\subset[\delta(\rho^{-},\alpha),\delta(\rho,\alpha))$ be a sequence converging to $\delta$.
@@
Case 2: Suppose $a(\delta,\rho^{+},\alpha)\geq 1$. As $a\in[0,1]$ satisfies (12), we have $a(\delta,\rho^{+},\alpha)=1=a$. If $\delta\in[\delta(\rho^{-},\alpha),\delta(\rho,\alpha))$, then $\phi(1^{-})=\phi(0^{-})+1<1$ by Lemma 4.2, which implies $\phi(\alpha^{-})\leq\phi(1^{-})<1$ by Proposition 3.6 and $a(\delta,\rho^{+},\alpha)<1$ by Lemma 4.3, which is a contradiction. It follows that $\delta=\delta(\rho,\alpha)$ and $\delta_{n}<\delta$ for all $n\in\mathbb{N}$. As $\delta^{\prime}\mapsto a(\delta^{\prime},\rho^{+},\alpha)$ is increasing and continuous, there exists $n_{0}\in\mathbb{N}$ such that $10<a(\delta_{n},\rho^{+},\alpha)<a(\delta,\rho^{+},\alpha)=$ 1 for all $n\geq n_{0}$. Let $(a_{n})_{n\in\mathbb{N}}$ be such that $a_{n}\in(a(\delta_{n},\rho^{+},\alpha),a(\delta_{n},\rho,\alpha)]\cap[0,1]$ for every $n\geq n_{0}$. Then, $(a_{n})_{n\in\mathbb{N}}$ converges to $a$, since $a(\delta_{n},\rho^{+},\alpha)$ converges to $a(\delta,\rho^{+},\alpha)=a=1$. As $(\delta_{n},a_{n})$ satisfies (44) for all $n\geq n_{0}$, we have $\rho(F_{\delta_{n},a_{n}})=\rho$ for all $n\geq n_{0}$ and $\rho(F_{\delta,a})=\rho$, by Lemma A.3.
@@
We start with studying the position of interval $[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]$ according to the position of $\alpha$ with respect to $1-\rho$.
@@
For any $\rho\in(0,1)$ we have $[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda-d,1-d)$ if $\alpha\in[0,1-\rho)$ and $[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda,1)$ if $\alpha\in(1-\rho,1]$.

### v3

@@ 2.2 Main results
For any symbolic sequence $i\in\{0,1,2\}^{\mathbb{N}}$, the complexity function $p_{i}:\mathbb{N}\setminus\{0\}\to\mathbb{N}$ of $i$ is defined by $p_{i}(n)=\#\{i_{t}i_{t+1}\dots i_{t+n-1}:t\geq 0\}$ for every $n\geq 1$. That is, $p_{i}(n)$ is the number of distinct blocks of size $n$ of consecutive symbols appearing in $i$. The complexity of a periodic symbolic sequence is eventually equal to its period. The complexity of the codes of irrational rotations with respect to any given partition can be computed thanks to Theorem 10 of [1]. Applying this theorem, we obtain the following corollary of Theorem 2.10:
Corollary 2.13.
Let $\lambda\in(0,1)$, $d\in(0,1-\lambda)$, $\rho\in(0,1)\setminus\mathbb{Q}$ and $\alpha\in[0,1]$. Let $\delta\in(1-\lambda-d,1)$ and $a\in[0,1]$ satisfying (24) and denote $f_{\delta,a}$ the map (6). Let $\eta=\eta_{1}$ if $\alpha\in[0,1-\rho)$, $\eta=a$ if $\alpha=1-\rho$ and $\eta=\eta_{2}$ if $\alpha\in(1-\rho,1]$ and let $P$ be the partition of $[0,1)$ in two or three left closed intervals defined by the positions of $a$ and $\eta$ in $[0,1)$. Let $x\in\Lambda$ and $p:\mathbb{N}\setminus\{0\}\to\mathbb{N}$ be the complexity function of the code of the orbit of $x$ with respect to $P$. Then,
  1. 1.
    $P$ is the partition given by the positions of the discontinuities of $f_{\delta,a}$ in $[0,1)$.
  2. 2.
    If $\{\alpha\}\neq R_{\rho}^{k}(0)$ for every $k\in\mathbb{Z}$, we have
    $p(n)=2n+1\qquad\forall n\geq 1.$
  3. 3.
    If $\{\alpha\}=R_{\rho}^{k}(0)$ for some $k\in\mathbb{Z}$, we have
    $p(n)=n+b\qquad\forall n\geq\max\{1,b-1\},$
    where $b=|k|$ if $k\leq-2$, $b=1$ if $k\in\{-1,0\}$ and $b=k+1$ if $k\geq 1$.
Proof.
Suppose that $\delta\in(1-\lambda-d,1)$ and $a\in[0,1]$ satisfy (24). As $\rho\in(0,1)\setminus\mathbb{Q}$, by Proposition 2.4, the map $f_{\delta,a}$ is of the type (9) if $\alpha\in[0,1-\rho)$, of the type (11) if $\alpha=1-\rho$ and of the type (10) if $\alpha\in(1-\rho,1]$. Therefore, $P$ is the partition given by the positions of the discontinuities of $f_{\delta,a}$ in $[0,1)$. Now, if $x\in\Lambda$, then $x=\phi_{\delta,\rho,\alpha}(y)$ for some $y\in[0,1)$, since $\phi_{\delta,\rho,\alpha}([0,1))=\Lambda$ for $\rho\in(0,1)\setminus\mathbb{Q}$. Then by item 3) of Theorem 2.10, the code of the orbit of $x$ by $f_{\delta,a}$ with respect to $P$ is the same as the code of the orbit of $y$ by $R_{\rho}$ with respect to the partition of $[0,1)$ in two or three left closed intervals defined by the positions of $\alpha$ and $1-\rho$ in $[0,1)$. Applying Theorem 10 of [1], we obtain that the complexity function of such a rotation’s code is given by item 2) or 3). ∎
Example 2.14.
Here we give an example of application of Corollary 2.13. Choose any $\lambda\in(0,1)$, $d\in(0,1-\lambda)$ and let $\rho=\pi/4$. Now, by choosing a suitable $\alpha$, we can obtain maps whose symbolic dynamics has a desired admissible complexity:
  • Let $\alpha=1/4\neq R^{k}_{\frac{\pi}{4}}(0)$ for every $k\in\mathbb{Z}$. Let $\delta=\delta(\pi/4,1/4)$ and $a=a(\delta,\pi/4,1/4)$. Then, $f_{\delta,a}$ is of the type (10) since $\alpha=1/4>1-\rho=(4-\pi)/4$. By items 1) and 2) of Corollary 2.13, for any $x\in\Lambda$ the code of the orbit of $x$, with respect to the partition $P_{\eta_{2},a}:=\{[0,\eta_{2}),[\eta_{2},a),[a,1)\}$, has complexity $p(n)=2n+1$ for all $n\geq 1$.
  • Let $\alpha=R^{k}_{\frac{\pi}{4}}(0)$ for $k=100$, that is $\alpha=\{25\pi\}$. Let $\delta=\delta(\pi/4,\{25\pi\})$ and
    $a\in[a(\delta,(\pi/4)^{+},\{25\pi\}),a(\delta,\pi/4,\{25\pi\})]=[a(\delta,\pi/4,\{25\pi\})-\lambda^{99}(1-\lambda-d),a(\delta,\pi/4,\{25\pi\})].$
    Then, $f_{\delta,a}$ is of the type (10) since $\alpha=\{25\pi\}>1-\rho=(4-\pi)/4$. By items 1) and 3) of Corollary 2.13, for any $x\in\Lambda$ the code of the orbit of $x$, with respect to the partition $P_{\eta_{2},a}$, has complexity $p(n)=n+101$ for all $n\geq 100$.
@@
The function $\phi_{\delta,\rho,\alpha}$ allows to parametrize a large variety of Cantor sets. Indeed, choose any $\rho\in(0,1)\setminus\mathbb{Q}$ any $\alpha\in[0,1]$. Then the closure of $\phi_{\delta,\rho,\alpha}([0,1))$ is a Cantor set, whose extremities are 0 and $1$ if $\delta=\delta(\rho,\alpha)$ and $\{\alpha+k\rho\}\neq 0$ for all $k\geq 1$, see Lemma 6.2. In such a case $\phi_{\delta,\rho,\alpha}(y)$ can be written as
$\phi_{\delta,\rho,\alpha}(y)=\phi_{\delta(\rho,1),\rho,1}(y)+\frac{d}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}(\theta_{\alpha}(\{y-k\rho\})-\theta_{\alpha}(\{-k\rho\}))\qquad\forall y\in[0,1).$ (28)
The closure of $\phi_{\delta(\rho,1),\rho,1}([0,1))$ is the Cantor set $C_{1}$ corresponding to the attractor of the map $x\mapsto\lambda x+\delta\pmod{1}$. In [7], it has been shown that $C_{1}\setminus\{0,1\}$ contains only transcendental numbers if $\lambda=1/b$, where $b\geq 2$ is an integer. It would be interesting to study such a property for the Cantor set obtained from (28) as a function of the values of $d\in(0,1-\lambda)$ and $\alpha\in(0,1)$.
In the sequel of the paper we prove the results presented above. As mentioned before, we always suppose $\lambda\in(0,1)$ and $d\in(0,1-\lambda)$, as stated in the definition (4) of $F$, and do not recall it from now on. Also, we will mostly write $\phi$ for $\phi_{\delta,\rho,\alpha}$ and $f$ for $f_{\delta,a}$.
@@ 3 Construction and general properties of $\phi$
is the lift of the rotation of angle $\rho$. In this section, we first show why the function $\phi=\phi_{\delta,\rho,\alpha}$ given in (23) is a natural candidate and then we establish conditions on this function which ensure (29).
@@ 3.3 Monotonicity properties of $\phi$
We have proved that $\psi_{\alpha}$ is non-decreasing and item $1)$. Finally, if $z\notin\mathbb{Z}\cup(\mathbb{Z}+\alpha)$ then, the integer part and the fractional part are continuous at $z$ and $\theta_{\alpha}$ is continuous at $\{z\}$, which implies that $\psi_{\alpha}(z)=\psi_{\alpha}(z^{-})$. Suppose $z\in\mathbb{Z}\cup(\mathbb{Z}+\alpha)$. If $\alpha\in(0,1)$, then either $z\in\mathbb{Z}$ and $\psi_{\alpha}(z)-\psi_{\alpha}(z^{-})=1-\lambda-d$ or $z\in\mathbb{Z}+\alpha$ and $\psi_{\alpha}(z)-\psi_{\alpha}(z^{-})=d$. If $\alpha\in\{0,1\}$, then $z\in\mathbb{Z}\cap(\mathbb{Z}+\alpha)$ and $\psi_{\alpha}(z)-\psi_{\alpha}(z^{-})=1-\lambda$. This ends the proof of (35). ∎
@@
The continuity properties of $\phi$ follow from those of $\psi_{\alpha}$ stated in Lemma 3.5. Now, if $y^{\prime}>y\in\mathbb{R}$, then by $1)$ of Lemma 3.5
@@
which proves that $\phi$ is a non-decreasing function. Also, by (36), $\phi(y^{\prime})-\phi(y)=0$ if and only if $\psi_{\alpha}(y^{\prime}-k\rho)-\psi_{\alpha}(y-k\rho)=0$ for every $k\geq 1$. By $1)$ of Lemma 3.5, this can hold only if
@@
If $\rho\in(0,1)\setminus\mathbb{Q}$, then the sequence $(\{y^{\prime}-k\rho\})_{k\in\mathbb{N}}$ is dense in $[0,1]$. Therefore, there exists $k\in\mathbb{N}$ such that $0\leq\{y^{\prime}-k\rho\}<y^{\prime}-y=y^{\prime}-k\rho-(y-k\rho)$ and
@@
Let $\alpha\in[0,1]$. If $\rho\in(0,1)\cap\mathbb{Q}$, then $\phi([0,1))\subset[0,1)$ if and only if $0\in(\phi(0^{-}),\phi(0)]$. If $\rho\in(0,1)\setminus\mathbb{Q}$, then the same holds replacing $(\phi(0^{-}),\phi(0)]$ with $\phi(0^{-}),\phi(0)]$.
@@
For $\rho\in(0,1)\setminus\mathbb{Q}$, 1) and 2) hold replacing $(\phi(\alpha^{-}),\phi(\alpha)]$ with $\phi(\alpha^{-}),\phi(\alpha)]$.
@@
$1)$ Let $\alpha\in(0,1)$. Suppose that $\theta_{\alpha}(y)=\theta_{a}(\phi(y))$ for all $y\in[0,1)$. So, for every $y\in[0,1)$ we have, $\phi(y)<a$ if and only $y<\alpha$. This implies that $\phi(\alpha^{-})\leq a\leq\phi(\alpha)$. In the case $\rho\in(0,1)\cap\mathbb{Q}$, by Lemma 3.8, there exists $y\in[0,\alpha)$ such that $\phi(y)=\phi(\alpha^{-})$ and therefore $\phi(\alpha^{-})<a$.
@@
$2)$ Let $\alpha=0$. If $\phi([0,1))\subset[0,1)$, by Lemma 3.9 we have $\phi(\alpha^{-})=\phi(0^{-})<0\leq a$ if $\rho\in(0,1)\cap\mathbb{Q}$ and $\phi(\alpha^{-})=\phi(0^{-})\leq 0\leq a$ if $\rho\in(0,1)\setminus\mathbb{Q}$. Suppose $1=\theta_{0}(y)=\theta_{a}(\phi(y))$ for every $y\in[0,1)$. Then, $\phi(\alpha)=\phi(0)\geq a$ and it follows that $a\in[\phi(\alpha^{-}),\phi(\alpha)]$ with left open interval if $\rho\in(0,1)\cap\mathbb{Q}$. Conversely, if $a\leq\phi(\alpha)=\phi(0)$, then, by monotonicity of $\phi$, $a\leq\phi(y)$ and $1=\theta_{a}(\phi(y))=\theta_{0}(y)$ for every $y\in[0,1)$.
@@
$=\frac{1}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}(\psi_{\alpha}(\alpha-k\rho)-\psi_{\alpha}((\alpha-k\rho)^{-}))$
$=\frac{1}{\lambda}\sum_{k=1}^{\infty}\lambda^{k}\left((1-\lambda-d)\mathbb{1}_{\mathbb{Z}}(\alpha-k\rho)+d\mathbb{1}_{\mathbb{Z}+\alpha}(\alpha-k\rho)\right)$
@@
The following two lemmata give the the monotonicity properties of $\delta(\rho,\alpha)$ with $\rho$ and allow to show item 1) of Theorem 2.3.
Lemma 4.5.
Also, for any $\rho\in(0,1)$ we have $\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda-d,1)$.
Proof.
As $\psi_{1-\alpha}$ is right continuous, we have
On the other hand,
Lemma 4.6.
Let $\rho\in(0,1)$, $\alpha\in[0,1]$ and suppose $\delta\in[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]$. Then, $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]\neq\emptyset$ and $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,1]$ if $\alpha\in(0,1)$ or $\rho\notin\mathbb{Q}$.
Proof.
As $\delta\in[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]$, we deduce from (42) that $\phi(0^{-})\leq 0\leq\phi(0)$, which together with (32) implies $\phi(1^{-})\leq 1\leq\phi(1)$. Since $\phi$ is non-decreasing, we have $0\leq\phi(\alpha)$ and $\phi(\alpha^{-})\leq 1$ for every $\alpha\in[0,1]$. From (44) we obtain that $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]\neq\emptyset$. Now, if $\alpha\in(0,1)$ then $0\leq\phi(\alpha^{-})\leq\phi(\alpha)\leq 1$. If $\alpha\in\{0,1\}$ and $\rho\notin\mathbb{Q}$, then $\phi$ is continuous in $\{0,1\}$ and $\phi(\alpha^{-})=\phi(\alpha)\in\{0,1\}$. We conclude from (42) that $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,1]$ if $\alpha\in(0,1)$ or $\rho\notin\mathbb{Q}$. ∎
@@
First, $rp/q\in\mathbb{Z}$ only if $r=0$, since $p$ and $q$ are co-prime. Second, the set $D_{\rho,\alpha}$ is either empty or contains a unique element $r_{\alpha}$. Indeed, if there exist $r_{0}$ and $r_{1}$ in this set, then $|r_{1}-r_{0}|/q\in\mathbb{Z}$ because $p$ and $q$ are co-prime, and it follows that $r_{0}=r_{1}$ since $0\leq|r_{1}-r_{0}|/q<1$. Also, the set $\{r\in\{1,\dots,q-1\}:-rp/q\in\mathbb{Z}+\alpha\}$ is empty if and only if $D_{\rho,\alpha}$ is empty and, if it is not empty, then it contains a unique element $r^{\prime}_{\alpha}=-(r_{\alpha}-q)$. It follows that,
@@
Now, we give in the following lemma the expressions of $\delta(\rho,\alpha)$ and $a(\delta,\rho,\alpha)$ for $\rho\in\mathbb{Q}$.
@@
Let $\rho\in(0,1)$ and $\alpha\in[0,1]$. The item 1) of Theorem 2.3 is deduced from Lemma 4.5 and Lemma 4.6. The items 2) and 3) follow from Lemma 4.7 and Lemma 4.8. Now, suppose that $\delta\in(1-\lambda-d,1)$ and $a\in[0,1]$ satisfy
@@
Also, $\phi([0,1))\subset[0,1)$ by Proposition 3.4, which implies that $\left\lfloor\phi(y)\right\rfloor=\left\lfloor y\right\rfloor$ for all $y\in\mathbb{R}$ by $2)$ of Lemma 3.3. Let $x=\phi(0)$, then
@@
To end the proof of Theorem 2.3, let us suppose that $\rho\in\mathbb{Q}$ and that $(\delta,a)$ satisfies (14) with $\delta=\delta(\rho,\alpha)$ or $a=a(\delta,\rho^{+},\alpha)$. Using continuity arguments of the rotation number given in [29] and adapted to our two parameters context in Appendix A.1, let us show that the rotation number of $F$ is $\rho$. As $\rho$ is a rational number, by Lemma 4.8, the intervals $\delta(\rho^{-},\alpha),\delta(\rho,\alpha))$ and $(a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]$ are non-empty. So, let $(\delta_{n})_{n\in\mathbb{N}}\subset[\delta(\rho^{-},\alpha),\delta(\rho,\alpha))$ be a sequence converging to $\delta$.
@@
Case 2: Suppose $a(\delta,\rho^{+},\alpha)\geq 1$. As $a\in[0,1]$ satisfies (14), we have $a(\delta,\rho^{+},\alpha)=1=a$. If $\delta\in[\delta(\rho^{-},\alpha),\delta(\rho,\alpha))$, then $\phi(1^{-})=\phi(0^{-})+1<1$ by Lemma 4.2, which implies $\phi(\alpha^{-})\leq\phi(1^{-})<1$ by Proposition 3.6 and $a(\delta,\rho^{+},\alpha)<1$ by Lemma 4.3, which is a contradiction. It follows that $\delta=\delta(\rho,\alpha)$ and $\delta_{n}<\delta$ for all $n\in\mathbb{N}$. As $\delta^{\prime}\mapsto a(\delta^{\prime},\rho^{+},\alpha)$ is increasing and continuous, there exists $n_{0}\in\mathbb{N}$ such that $0<a(\delta_{n},\rho^{+},\alpha)<a(\delta,\rho^{+},\alpha)=1$ for all $n\geq n_{0}$. Let $(a_{n})_{n\in\mathbb{N}}$ be such that $a_{n}\in(a(\delta_{n},\rho^{+},\alpha),a(\delta_{n},\rho,\alpha)]\cap[0,1]$ for every $n\geq n_{0}$. Then, $(a_{n})_{n\in\mathbb{N}}$ converges to $a$, since $a(\delta_{n},\rho^{+},\alpha)$ converges to $a(\delta,\rho^{+},\alpha)=a=1$. As $(\delta_{n},a_{n})$ satisfies (48) for all $n\geq n_{0}$, we have $\rho(F_{\delta_{n},a_{n}})=\rho$ for all $n\geq n_{0}$ and $\rho(F_{\delta,a})=\rho$, by Lemma A.3.
@@
We start with studying the position of interval $\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]$ according to the position of $\alpha$ with respect to $1-\rho$.
@@
For any $\rho\in(0,1)$ we have $\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda-d,1-d)$ if $\alpha\in[0,1-\rho)$ and $\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda,1)$ if $\alpha\in(1-\rho,1]$.